Suppose an investor is given an investment opportunity that will provide a future value of Kshs 50,000 when it matures at the end of two years. How much would an earning of Kshs 50,000 be at the end of 2 years, at a rate of 10 percent?


1
Expert's answer
2022-02-15T04:20:56-0500

FV= PV(1+r)"^n"

"PV= \frac{FV}{(1+r)^n}"


"\frac{50000}{(1.1)^2}= \frac{50000}{1.21}"


= 41,322